In a research study, you measured the weights of all learners in a college freshman class. The normally distributed set of scores for the population of learners has a mean of 120 pounds and a standard deviation of 8 pounds. (a) If you randomly sampled one score from the population, what is the probability it would be equal to or greater than a score of 134? (b) How would you briefly explain the result obtained to another person without your statistical knowledge regarding probability?
Answer by stanbon(75887)

In a research study, you measured the weights of all learners in a college freshman class. The normally distributed set of scores for the population of learners has a mean of 120 pounds and a standard deviation of 8 pounds.
------------------------------
(a) If you randomly sampled one score from the population, what is the probability it would be equal to or greater than a score of 134?
z(134) = (134-120)/8 = 1.75
P(x >= 134) = P(z >= 1.75) = 0.0401
-------------------------------------------------
(b) How would you briefly explain the result obtained to another person without your statistical knowledge regarding probability?
I'll leave that to you.
Cheers,
Stan H.
===========